OpenStudy (anonymous):

]If 85 g of NaCl are dissolved in 1.4 kg of water, what would be the expected change in boiling point? The boiling point constant for water (Kb) is 0.51 °C/m. 0.53°C 1.1°C 1.5°C 2.1°C

OpenStudy (anonymous):

@Ash90

OpenStudy (anonymous):

sorry about that my computer is lagging and i got shut off. but here is how you do this one but this can be tricky: Delta T = mKb , Kb, m = molality = moles solute/kg solvent = 85/58.5/1.4=1.03 1.03*0.51=0.53 c

OpenStudy (anonymous):

No you seriously rock and are like the most helpful person ever!! @Ash90

OpenStudy (anonymous):

Omg I'm so sorry but is there any way you could help me with one more?? @Ash90 I'll tag you in it!! <3

OpenStudy (anonymous):

let me know if i've answered them all :)